Suppose that the Sales in December 2019 were 46,595. The sales are expected to grow by 6% in January 2020, by 9% in February 2020, and by 6% in March 2020. After that, the sales are expected to grow by 10%.

The company's cash collection is as follows: 65% is collected in the same month and the remainder in the next month.

What is the cash collection for July, 2020?

Enter your answer rounded off to two decimal points. Do not enter $ or comma in the answer box.
Answer #1

The total Cash collection in July =79664.35

Attached is the calculationCash in Decembar ash un damlary und = 30286.75 - 32103.95 +16308.25 - 4.8412:2 = 349.92.75 + 16944.27 5193702 Cash in feb Cassales in december = 46595 January - 49.390°7 febrny = 53835.86 marcha 53835.867 3230.15 o 570 66. O11 April - 57066.0l1 + 570